Hi need help with this question, am in year 5.
COuld you please also show me the full working out.
Thank-you. i tried but couldn't get the answer.
linda ran 1/3 hour at 12km/h. She walked another 1/2 hour at 6km/h. Find the average speed of the whole whole journey.

Answer by htmentor(1343)

She ran a distance of 12 km/h x 1/3 h = 4 km
She walked for a distance of 6 km/h x 1/2 h = 3 km
Average speed = distance/time = (4 km + 3 km)/(1/3 h + 1/2 h) = 7/(5/6) = 42/5 = 8.4 km/h